Don, a canteen owner claims that the average meal cost of his usual customer is 190 pesos. In order to test his claim , don took a random sample of 25 customers and found out that the meal cost is 210 with a standard deviation of 30 pesos. Test the hypothesis at 0.01 level of significance

1
Expert's answer
2022-06-07T13:50:53-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=190"

"H_1:\mu\not=190"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=24" and the critical value for a two-tailed test is "t_c =2.79694."

The rejection region for this two-tailed test is "R = \{t:|t|>2.79694\}."


The t-statistic is computed as follows:



"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{210-190}{30/\sqrt{25}}=3.333333"


Since it is observed that "|t|=3.333333>2.79694=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for two-tailed, "df=24" degrees of freedom, "t=3.333333" is "p=0.002776," and since "p=0.002776<0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is different than 190, at the "\alpha = 0.01" significance level.